OpenStudy (anonymous):

How to Factor? x^3+4x^2-x-4

OpenStudy (anonymous):

\[x^3+4x^2-x-4\]

OpenStudy (mathmath333):

\(\large\tt \color{black}{x^3+4x^2-x-4}\) \(\large\tt \color{black}{ =x^2 (x+4)-(x+4) }\) \(\large\tt \color{black}{=(x+4)(x^2-1)}\) \(\large\tt \color{black}{=(x+4)(x-1)(x+1)}\)